### v1

# HOI-Dyn: Learning Interaction Dynamics for Human-Object Motion Diffusion

Abstract
Generating realistic 3D human-object interactions (HOIs) remains a challenging task due to the difficulty of modeling detailed interaction dynamics. Existing methods treat human and object motions independently, resulting in physically implausible and causally inconsistent behaviors. In this work, we present HOI-Dyn, a novel framework that formulates HOI generation as a driver-responder system, where human actions drive object responses. At the core of our method is a lightweight transformer-based interaction dynamics model that explicitly predicts how objects should react to human motion. To further enforce consistency, we introduce a residual-based dynamics loss that mitigates the impact of dynamics prediction errors and prevents misleading optimization signals. The dynamics model is used only during training, preserving inference efficiency. Through extensive qualitative and quantitative experiments, we demonstrate that our approach not only enhances the quality of HOI generation but also establishes a feasible metric for evaluating the quality of generated interactions.
1 Introduction
Synthesizing complex and realistic 3D human-object interactions (HOIs) is essential for progress in VR/AR, computer animation, and robotics [1, 2, 3, 4], yet remains a significant challenge. Compared to human motion generation—whether for single or multiple people [5, 6]—HOI is significantly more difficult. Human motion generation typically involves relatively free movement, making it easier to generate plausible sequences [7, 8]. However, HOI requires capturing intricate interaction dynamics, such as stable contact, forces, and action-response relationships. Simply applying human motion generation frameworks to HOI often produces independent motion of the human and object, leading to physically unrealistic and causally inconsistent behaviors.
Previous works on HOI generation mainly focused on interactions between humans and static objects or scenes [9, 10, 11, 12], such as sitting on a sofa. Recent advances have shifted toward controllable synthesis of dynamic HOI, where both the human and object move synchronously [13, 14]. For example, given an initial state and a textual instruction, the system can generate a motion sequence where a human picks up an object, such as a bench, and places it elsewhere. This approach enables more flexible motion generation and a wide range of applications.
However, existing methods often fail to capture the core interaction dynamics between humans and objects. These approaches typically focus on modeling either object affordances or contact points [1, 4, 15], or simply integrating human and object motions through diffusion-based models [2, 13]. However, they do not fully address how objects should respond to human actions, often leading to physical and causal inconsistencies.
In this work, we propose a new perspective: framing HOI generation as a driver-responder system [16], where human actions serve as the driver and objects respond accordingly. At the heart of this approach is the modeling of interaction dynamics, which describes how objects should naturally react to human motions. This view offers several advantages:
  • Contact is implicitly governed by the dynamics—no need to explicitly model it. If there is no contact, there is no response; if contact occurs, the object’s response is naturally determined by the interaction dynamics.
  • Object motion is not independent—each step of their movement is driven by the human’s actions and controlled through specific instructions or context, ensuring a coherent and physically plausible interaction.
Building on this perspective, we design a new HOI generation framework that explicitly incorporates interaction dynamics into the motion synthesis process, yielding state-of-the-art performance on challenging HOI benchmarks and offering a physically grounded solution to HOI generation. Specifically, our contributions are as follows:
  • We introduce a novel driver–responder formulation for HOI generation from a synchronized control perspective, modeling the causal dependencies between human actions and object responses in a dynamic and physically consistent manner.
  • We propose a lightweight transformer-based interaction dynamics model that answers how objects should react dynamically to human actions, taking into account the context of human motion and specific contact situations.
  • We introduce a residual-based interaction dynamics loss that serves HOI motion diffusion, compensating for prediction noise in the dynamics model. This loss helps prevent misleading optimization gradients and ensures the focus remains on core generative inconsistencies, thereby improving the quality of the generated motion sequences.
  • We demonstrate the effectiveness of our approach through extensive qualitative and quantitative experiments, highlighting that the proposed model not only improves HOI generation but also serves as a reliable evaluation metric for assessing the quality of generated interactions.
2 Related Work
2.1 Driver-Responder Synchronization
Driver-Responder Synchronization (master-slave synchronization) is observed in many coupled systems, such as those found in biological and physical domains [17, 18, 19]. These systems often rely on controllers to achieve synchronization, employing strategies such as adaptive and feedback control [20]. From a control theory perspective, synchronization is achieved by adjusting the coupling strength and applying feedback to align the system states [21]. This concept has been applied to various domains, but similar challenges arise in HOI generation.
Current HOI generation frameworks, however, often treat human and object motions as independent entities [13, 22, 23], which can lead to physically implausible object behaviors. Existing generative approaches [24] model the overall consistency between human and object, but they fail to account for the fine-grained interaction dynamics at the point of contact. To address this gap, we propose conceptualizing HOI generation as a Driver-Responder System, where human actions serve as the driver, controlling the object’s response. This perspective ensures a more coherent and physically consistent interaction by capturing the causal relationship between human movements and object reactions. We argue that this approach overcomes the limitations of previous models by introducing a refined internal control mechanism that improves the physical realism and causal consistency of the generated interactions.
2.2 Object and Contact Modeling
A substantial body of work on HOI generation has been dedicated to modeling object affordances and contact prediction. The method HOI-Diff [25] introduces an affordance prediction module to constrain interaction regions, by incorporating the estimated contact points into classifier-guidance to encourage close proximity between humans and objects. This method is further enhanced in CG-HOI [15] by explicitly predicting contact fields on the human body mesh and using them as strong priors during both training and inference. The method OMOMO [1] adopts a two-stage pipeline: it first infers contact points and then generates HOI sequences conditioned on these points. While these methods successfully model static interaction cues, they largely overlook the underlying interaction dynamics, i.e. how objects should physically and causally respond to human actions over time. In practice, accurate contact prediction is inherently difficult, and even with correct contact estimation, these approaches often fail to produce temporally coherent and physically realistic object behaviors.
2.3 Joint HOI Motion Generation
Another line of research explores joint generation of human-object motion using diffusion-based models. The model CHOIS [13] encodes human poses, object states, and limb-contact labels as part of the noising-denoising process. Their work enforces contact constraints by introducing additional object pose supervision during training and applying hard constraints during sampling. The method THOR [14] leverages a relationally-aware diffusion model that initiates human-object interaction guided by text, and intervenes in object motion based on learned human-object relations at each diffusion step.
While these methods integrate human and object motions within a unified generative framework and produce globally coherent sequences, they typically treat interaction as a high-level constraint or relational prior. Consequently, they often fail to model fine-grained interaction dynamics—specifically, how an object should physically and causally respond to human motion. This implicit treatment leads to plausible-looking results that may still violate physical consistency or lack local causal coherence.
3 Methodology
Our goal is to synthesize synchronized HOIs while maintaining internal causal consistency, by leveraging controllable signals such as textual descriptions and object geometry. We propose the HOI-Dyn framework (see Fig. 1) that explicitly models interaction dynamics, enabling the generation of more plausible and coherent motion sequences. The framework consists of two key components: Motion Diffusion and Interaction Dynamics. The Motion Diffusion component, based on a Transformer-based conditional diffusion model, jointly encodes the human, object, and interaction context into a unified representation. The Interaction Dynamics component provides auxiliary supervision to reinforce fine-grained causal consistency during motion generation.
Figure 1: Overview of the proposed HOI-Dyn framework. (a) Conditional Motion Diffusion synthesizes human-object interactions $\hat{\tau}_{0}=\{\hat{H},\hat{O},\hat{X}\}$ using a Transformer-based diffusion model, where $\hat{H}:=\{\hat{h}_{t}\}_{t=0}^{T-1}$ and $\hat{O}:=\{\hat{o}_{t}\}_{t=0}^{T-1}$. (b) The full framework integrates motion generation with interaction dynamics supervision. (c) Interaction Dynamics models object responses $\Delta\hat{o}_{t}^{*}$ based on human relative motion $\Delta\hat{h}_{t}$, object pose $\hat{o}_{t}$, and interaction context $\hat{s}_{t}$.
3.1 Interaction Dynamics
In diffusion-based motion generation, the output typically reflects clean motion rather than noise. The denoising process allows the model to internalize human-specific dynamics, resulting in physically plausible free movement. In contrast, object motion is externally driven and inherently constrained by physical laws—it cannot occur autonomously. To reflect this asymmetry, we decouple their roles as follows: the human follows internally guided dynamics, while the object responds to external control. This interaction is formalized as
[…]
where $h^{(t)}$ and $o^{(t)}$ denote the latent states of the human and the object at time step $t$, respectively. The functions $F_{h}(\cdot)$ and $F_{o}(\cdot)$ describe their respective internal dynamics, while $g_{h}$ and $g_{o}$ project these latent states to the observable outputs $y_{h}^{(t)}$ and $y_{o}^{(t)}$. $\Delta t$ is the sampling time step. The control signal $u^{(t)}$ is determined based on the error feedback $e(y_{h}^{(t)},y_{o}^{(t)})$, which reflects the discrepancy between the human’s intent and the object’s behavior. The term $s^{(t)}$ denotes the interaction context, which includes factors such as contact state, object geometry, and other environment- or task-specific conditions influencing the object’s motion.
In practice, we observe that existing diffusion models often fail to generate causally consistent interactions due to the absence of precise control signal $u^{(t)}$, which is essential for aligning the object’s motion with human intent [13, 14]. To address this limitation, we propose the interaction dynamics as a supervisory signal to implicitly optimize it in the diffusion process. Based on  (1), we derive the object’s relative motion as
$\Delta o^{(t)}=o^{(t+1)}-o^{(t)}=\Delta t\cdot F_{o}(o^{(t)},s^{(t)},u^{(t)})\approx\mathcal{D}(s^{(t)},o^{(t)},\Delta h^{(t)};\theta_{\mathcal{D}}),$ (2)
where $\Delta h^{(t)}=h^{(t+1)}-h^{(t)}$ denotes the human’s relative motion, and $\mathcal{D}(\cdot)$ is a learnable function parameterized by $\theta_{\mathcal{D}}$. This formulation emphasizes that object dynamics are governed not only by their internal state but also by human-induced interactions and the context.
To enhance sensitivity to varying interaction magnitudes, we extend the HOI prediction horizon from 1 to $k$, with $k$ being selected randomly from $[1,K]$. As shown in Fig. 1(c), the model takes as input the current object state $o^{(t)}$, interaction context $s^{(t)}$, and the cumulative human motion over the horizon $\Delta h_{t\to t+k}=h^{(t+k)}-h^{(t)}$, and predicts the corresponding object motion:
${\Delta o}_{t\to t+k}^{*}\approx\mathcal{D}(s^{(t)},o^{(t)},\Delta h_{t\to t+k};\theta_{\mathcal{D}}).$ (3)
The above predicted motion is further represented as a rigid-body transformation comprising a rotation ^$\hat{\mathcal{R}}^{(t\to t+k)}\in\mathrm{SO}(3)$ and translation $\hat{\mathcal{T}}^{(t\to t+k)}\in\mathbb{R}^{3}$, i.e., ${\Delta o}^{*}_{=}[\hat{\mathcal{R}}|\hat{\mathcal{T}}]$, applied to a set of object’s points $\mathcal{P}^{(t)}$ to yield the predicted future configuration as follows:
$\hat{\mathcal{P}}^{(t+k)}=\hat{\mathcal{R}}^{(t\to t+k)}\mathcal{P}^{(t)}+\hat{\mathcal{T}}^{(t\to t+k)}.$ (4)
To ensure that ^$\hat{\mathcal{R}}^{(t\to t+k)}$ is a valid rotation matrix, we apply singular value decomposition (SVD)-based projection to the raw network output ~$\tilde{\mathcal{R}}\in\mathbb{R}^{3\times 3}$, such that ~$\tilde{\mathcal{R}}=U\Sigma V^{\top}$ and ^$\hat{\mathcal{R}}=UV^{\top}$. We then define the following cost for object dynamics to quantify the error between the transformed keypoints:
$\Phi(\Delta o_{t\to t+k},{\Delta o}_{t\to t+k}^{*})=\|\mathcal{P}^{(t+k)}-\hat{\mathcal{P}}^{(t+k)}\|_{1}.$ (5)
The overall loss function is defined as the expected value over time steps $t$ and $k$ sampled from $\mathcal{U}(1,K)$ as follows:
$\mathcal{L}=\mathbb{E}_{t,\,k\sim\mathcal{U}(1,K)}\left[\frac{1}{k}\cdot\Phi(\Delta o_{t\to t+k},{\Delta o}^{*}_{t\to t+k})\right].$ (6)
This loss function guides the model to capture both the motion magnitude and the causal structure of interactions, yielding more realistic and consistent object motion in HOI.
3.2 Conditional Motion Diffusion
Having introduced the interaction dynamics, we now proceed to the motion generation stage. Here, we propose a conditional diffusion model that provides internal forces to synthesize temporally aligned motions for both the human and the object.
The joint human-object trajectory is denoted as $\tau=\{H,O,X\}$, with the human motion $H$, object motion $O$, and interaction context $X$ (e.g., hand-object and foot-object contact annotations). The context provides coarse guidance for synthesizing plausible interactions. We represent human motion using the SMPL-X parametric model [26], and represent each frame of the object motion using 3D translation and relative rotation. In the condition representation c, as shown in Fig. 1(a), we follow the CHOIS framework [13] to integrate contextual cues such as text prompts and the Basis Point Set (BPS) [27] of the object. More details can be found in Appendix B.
The complete conditional HOI Diffusion comprises both forward and reverse processes [28]. The forward process is modeled as a Markov chain over $N$ steps as follows:
$q(\tau_{1:N}|\tau_{0})=\prod_{n=1}^{N}q(\tau_{n}|\tau_{n-1}),\quad q(\tau_{n}|\tau_{n-1})=\mathcal{N}(\tau_{n};\sqrt{1-\beta_{n}}\,\tau_{n-1},\beta_{n}\mathbf{I}),$ (7)
where $\beta_{n}$ is a predefined noise schedule.
The reverse process is defined as
$p_{\theta}(\tau_{n-1}|\tau_{n},\mathbf{c})=\mathcal{N}(\tau_{n-1};\mu_{\theta}(\tau_{n},n,\mathbf{c}),\Sigma_{n}),$ (8)
where $\Sigma_{n}$ is a fixed variance, and $\mu_{\theta}=\frac{\sqrt{\alpha_{n}}(1-\bar{\alpha}_{n-1})}{1-\bar{\alpha}_{n}}\cdot\tau_{n}+\frac{\sqrt{\bar{\alpha}_{n-1}}\beta_{n}}{1-\bar{\alpha}_{n}}\cdot\hat{\tau}_{0}$, with $\bar{\alpha}_{n}=\prod_{i=1}^{n}\alpha_{i}$ and $\alpha_{n}=1-\beta_{n}$.
The reverse process (8) is modeled by a neural network $\theta_{\mathcal{G}}$ with the training loss:
$\mathcal{L}_{\text{hoi}}=\mathbb{E}_{\tau_{0},n}\left[\left\|\hat{\tau}_{0}(\tau_{n},n,\mathbf{c};\theta_{\mathcal{G}})-\tau_{0}\right\|_{1}\right],$ (9)
which minimizes the reconstruction error between the predicted and ground-truth trajectories.
Although the standard loss function (9) yields plausible motion sequences, physical inconsistencies—particularly in object trajectories—persist. To address this, we introduce an auxiliary Interaction Dynamics Loss based on a shared dynamics model $\mathcal{D}$. This loss penalizes discrepancies in object motion between generated and ground-truth sequences as follows:
$\mathcal{L}_{\text{dyn}}=\mathbb{E}_{t}\left[\left\|\Phi(\Delta\hat{o}_{t}^{*},\Delta\hat{o}_{t})-\Phi(\Delta o_{t}^{*},\Delta o_{t})\right\|_{1}\right],$ (10)
where $\{\Delta\hat{o}_{t}^{*}\}_{t=1}^{T-1}=\mathcal{D}(\Delta\hat{\tau}_{0})$ and $\{\Delta o_{t}^{*}\}_{t=1}^{T-1}=\mathcal{D}(\Delta\tau_{0})$ are the object motions predicted by $\mathcal{D}$ given the relative motion from the generated and ground-truth HOI $\hat{\tau}_{0}$ and $\tau_{0}$, respectively. The function $\Phi(\cdot,\cdot)$ computes the motion errors as described in (5). The per-frame residuals in $\mathcal{L}_{\text{dyn}}$ are
$\delta_{D,\text{gen}}=\Phi(\Delta\hat{o}_{t}^{*},\Delta\hat{o}_{t}),\quad\delta_{D,\text{gt}}=\Phi(\Delta o_{t}^{*},\Delta o_{t}).$ (11)
Ideally, if $\mathcal{D}$ perfectly captures the interaction dynamics, then $\delta_{D,\text{gt}}=0$ and the loss reverts to the direct supervision of $\delta_{D,\text{gen}}$. However, in practice, prediction errors are unavoidable due to imperfections in $\mathcal{D}$ and inaccuracies in the training data.
To mitigate the influence of these errors, we assume that $\mathcal{D}$ is locally smooth and time-homogeneous—its bias depends on the state but not on time $t$. This is reasonable since $\mathcal{D}$ predicts object dynamics from current states without explicit time dependence, and local smoothness reflects the continuity of physical motion. Under this assumption, when the generative and ground-truth trajectories exhibit similar state distributions, the error residuals tend to cancel out in expectation.
$\mathbb{E}_{t}[\delta_{D,\text{gen}}-\delta_{D,\text{gt}}]\approx 0.$ (12)
This residual formulation provides a robust supervisory signal: even if $\mathcal{D}$ is imperfect, its bias is removed through subtraction, allowing the learning algorithm to focus on the true inconsistencies in the generated motion. For a more detailed discussion, see Appendix C.
With this loss in place, each step of the generation process is guided to yield a more accurate control signal $u_{t}$, resulting in physically grounded and temporally coherent HOI sequences. We therefore extend the standard HOI training objective $\mathcal{L}_{\text{hoi}}$ (9) by incorporating both $\mathcal{L}_{\text{dyn}}$ (10) and an object-level reconstruction loss $\mathcal{L}_{\text{obj}}=\mathbb{E}_{t}[\Phi(o_{t},\hat{o}_{t})]$. All these losses are equally weighted to simplify the overall training process and avoid the need for complex hyperparameter tuning.
4 Experiments
Dataset.
We train and evaluate HOI generation using two datasets: (i) FullBodyManipulation, which provides 10 hours of high-quality paired object and human motion data involving 15 different objects [1]. We apply the OMOMO partitioning to this dataset: 15 subjects for training and 2 for testing. (ii) 3D-FUTURE, which consists of 3D models of various furniture items [29]. To assess the generalization capability, we select 17 objects and pair them with motion sequences from the FullBodyManipulation testing set.
Metrics.
We evaluate HOI from four aspects, as defined in CHOIS [13]: (i) Condition Matching, which measures the alignment of the predicted object trajectory with input waypoints using Euclidean distance at the start, end, and intermediate points (cm). (ii) Human Motion Quality, which is evaluated with the foot sliding score (FS), foot height $H_{\text{feet}}$, and Fréchet Inception Distance (FID) [30] that captures distributional differences between the generated and real motions. (iii) Interaction Quality: Based on the frame-wise contact labels, we compute the contact percentage $C_{\%}$, F1 score $C_{F_{1}}$, and hand-object penetration score $P_{\text{hand}}$ to assess physical plausibility. (iv) Ground Truth Difference, which includes the mean per-joint position error (MPJPE), root translation error $T_{\text{root}}$, object position error $T_{\text{obj}}$, and orientation error $R_{\text{obj}}$.
Implementation Details.
We train the interaction dynamics using FullBodyManipulation with $K=2$ as the maximum prediction horizon. The network has 0.5M parameters. We use the Adam optimizer with a learning rate of $1\times 10^{-3}$, adjusted via CosineAnnealingWarmRestarts [31] for 150 epochs with a batch size of 32. The model is then transferred to the HOI motion diffusion task, where fine-grained driver-responder relationships are encouraged during denoising, as in (10). Training is done from scratch with a learning rate of $1\times 10^{-4}$, batch size 32, and 100,000 steps. All experiments are conducted on a single NVIDIA RTX A4500 GPU, with total training time around 10 hours.
Baselines.
We compare our method HOI-Dyn with the state-of-the-art (SOTA) approach, CHOIS [13], whose results are reproduced under identical training conditions, including the same number of training steps. We also consider other baselines including the adapted versions of InterDiff [32], MDM [33], and OMOMO [1], whose results are borrowed from [13] for comparison. More specifically, InterDiff is modified to accept additional inputs, including text and sparse waypoints. MDM is updated to incorporate our object geometry representation and sparse waypoints, extending it to predict object motion. We introduce three variants of OMOMO: (i) Pred-OMOMO, which combines our object motion module with OMOMO; (ii) GT-OMOMO, which uses the ground truth object motion as input; and (iii) Lin-OMOMO, which incorporates a linear interpolation strategy to generate object motion trajectories and ensure consistent object rotation throughout the sequence.
4.1 Qualitative Results
We compare HOI-Dyn with the SOTA HOI model CHOIS from two complementary perspectives in a synchronized manner: (i) Action–Interaction–Response, which evaluates the physical and semantic plausibility of object motion before and after human actions; and (ii) Sequence-level Alignment, which assesses the consistency of the generated HOI sequence with the input conditions and its overall causal coherence. Together, these perspectives capture both the micro-level physical causality and macro-level semantic consistency.
As shown in Figs. 2 (a) and (b), CHOIS, which lacks interaction dynamics modeling, often causes premature object motion. The objects tend to move toward expected contact points (e.g., the hand) spontaneously before human actions begin, leading to implausible behaviors such as bouncing, wobbling, or sliding. In contrast, HOI-Dyn reduces such artifacts, promoting more natural human-object interactions. Fig. 2 (c) further highlights the difference in object responses after contact: HOI-Dyn generates subtle, physically plausible reactions (with slight sliding after a gentle kick), while CHOIS often produces exaggerated effects like flying or floating.
Fig. 2 (d) shows the keyframes for sequence-level comparison. Both methods generally respect input constraints, demonstrating the effectiveness of motion diffusion. However, CHOIS often exhibits disjoint human-object motion and weak causality, while HOI-Dyn ensures consistent object motion and maintains contact, resulting in more coherent and plausible HOI sequences. More visualization results can be found in AppendixD.
Figure 2: Comparison of HOI-Dyn and CHOIS on physical plausibility and sequence-level coherence. (a–b) CHOIS produces premature object motion lacking causal timing; (c) HOI-Dyn generates more realistic post-contact responses; (d) HOI-Dyn maintains consistent human-object interaction across the full sequence. Green markers indicate object initial state and sparse waypoints.
4.2 Quantitative Results
As shown in Table 1, HOI-Dyn consistently outperforms existing methods across all evaluation metrics. It achieves the lowest $T_{s}$ and $T_{e}$ in condition matching, indicating better alignment with input constraints. For human motion, it yields the best foot sliding score (FS) and FID, demonstrating a more realistic motion synthesis. In terms of interaction quality, HOI-Dyn attains the highest $C_{F1}$ and C%, while maintaining a comparable hand penetration score $P_{\text{hand}}$. It also achieves the lowest MPJPE, root translation $T_{\text{root}}$, and object motion errors ($T_{\text{obj}},R_{\text{obj}}$), demonstrating improved accuracy in modeling human-object dynamics.
Table 1: Comparison of methods across different metrics. Arrows indicate whether lower ($\downarrow$) or higher ($\uparrow$) is better, and the same notation applies hereafter.
Method Condition Matching Human Motion Interaction GT Difference
[…]
Interdiff 0.00 158.84 72.72 0.90 0.42 208.0 0.33 0.27 0.55 25.91 63.44 88.35 1.65
MDM 5.18 33.07 19.42 6.72 0.48 6.16 0.53 0.43 0.66 17.86 34.16 24.46 1.85
Lin-OMOMO 0.00 0.00 0.00 7.21 0.41 15.33 0.57 0.54 0.51 21.73 36.62 17.12 1.21
Pred-OMOMO 2.39 8.03 4.15 7.08 0.40 4.19 0.66 0.62 0.58 18.66 28.39 16.36 1.05
GT-OMOMO 0.00 0.00 0.00 7.10 0.41 5.69 0.67 0.59 0.55 15.82 24.75 0.00 0.00
CHOIS 2.10 6.16 3.03 3.39 0.41 0.87 0.66 0.54 0.61 16.01 24.33 14.29 0.99
HOI-Dyn (Ours) 1.75 5.58 3.26 3.07 0.37 0.48 0.71 0.60 0.64 15.60 23.90 12.47 0.90
To assess the generalizability of our model to novel objects, we also perform experiments on the 3D-FUTURE dataset. As shown in Table 2, although not trained on this dataset, our model significantly improves the quality of human motion and, more importantly, enhances the fidelity of human-object contact, with only a slight increase in penetration.
Table 2: Interaction synthesis results on the 3D-FUTURE dataset [29].
Method Condition Matching Human Motion Interaction
[…]
InterDiff 0.00 161.26 72.77 -0.26 0.42 207.3 0.24 0.11
MDM 12.58 40.55 28.72 7.02 0.49 8.50 0.34 0.26
Lin-OMOMO 0.00 0.00 0.00 6.32 0.42 23.17 0.44 0.11
Pred-OMOMO 4.15 9.03 3.89 6.08 0.40 3.74 0.50 0.18
CHOIS 3.23 6.21 2.99 2.95 0.42 1.67 0.47 0.19
HOI-Dyn (Ours) 4.60 6.17 2.95 2.56 0.37 1.62 0.54 0.26
4.3 Further Discussion
Effect of Interaction Dynamics Variants.
We analyze the effects of prediction horizon $K$, network design, and model complexity on the interaction dynamics performance. As shown in Fig. 4, a small $K$ limits large motion capture, while a large $K$ weakens subtle interaction modeling. Empirically, $K{=}2$ or $K{=}3$ yield the best results. For network design, we compare the decoupled and coupled motion strategies. Inspired by [34, 35, 36], the decoupled approach predicts object rotation and translation separately. Our coupled design instead models the unified influence of human motion and contact, yielding a better performance under similar parameter and FLOP constraints, highlighting the inherent coupling in HOI, as shown in Table 3. We also evaluate the model efficiency by varying the Transformer depth ($D$), feature dimension ($F$), and head count ($H$). The results in Table 3 show that a lightweight model with 0.5M parameters and 0.2 GFLOPs suffices to capture high-quality interaction dynamics. For a more detailed discussion, see Appendix E
Effect of Different Guidance.
Classifier-based guidance is often employed during inference in generative models. Following [13], we apply two types of guidance: feet-floor and hand-object. For a more detailed description, see Appendix F. As shown in Table 4, even without guidance, our method surpasses CHOIS, especially in the interaction quality. The feet-floor term enhances foot realism and reduces FID, while the hand-object term boosts contact accuracy but slightly harms motion quality. Combining both yields consistent gains across all metrics, achieving a new SOTA.
Dynamics as Causality Metric.
We evaluate the interaction dynamics modeling as a proxy for causal consistency. Specifically, we compute the point cloud loss $\mathbb{E}_{t}[\Phi(\Delta\hat{o}_{t}^{*},\Delta\hat{o}_{t})]$ in (5), where $\Delta\hat{o}_{t}$ is the object’s observed motion and $\Delta\hat{o}_{t}^{*}$ is the predicted response from the dynamics model. As shown in Fig. 4, despite some noise, HOI-Dyn closely aligns with the ground truth, while CHOIS shows a notable deviation, indicating a better causal modeling by HOI-Dyn. Furthermore, we visualize the frame-wise dynamics loss in Fig. 5, where the loss spikes when objects behave unrealistically. The results show that the loss remains low for HOI-Dyn, indicating that our model more effectively captures physically consistent interactions. These findings support the use of dynamics-based losses as a metric for causal evaluation in HOI synthesis.
Figure 3: Effect of Horizon $K$.
Figure 4: Object Loss via Dynamics.
Table 3: Effect of Design Variants. Our lightweight coupled model captures high-quality interaction dynamics, outperforming the decoupled variants under similar constraints.
Architecture Network Configuration Object Point Cloud Loss Params Flops
$K$=1 $K$=2 $K$=3 $K$=4 $K$=5 $K=10$ (M) (G)
Coupled (K=1) 0.316±0.0004 0.514±0.0012 0.763±0.0009 1.080±0.0046 1.444±0.0051 3.422±0.0050
Coupled (K=2) 0.313±0.0003 0.462±0.0003 0.622±0.0007 0.791±0.0024 0.982±0.0036 2.279±0.0059
Coupled (K=3) D4-F64-H8 0.313±0.0006 0.458±0.0016 0.603±0.0010 0.756±0.0024 0.920±0.0027 1.900±0.0079 0.483 0.201
Coupled (K=5) 0.321±0.0005 0.459±0.0012 0.599±0.0013 0.736±0.0011 0.879±0.0008 1.633±0.0031
Coupled (K=10) 0.329±0.0003 0.467±0.0022 0.604±0.0022 0.737±0.0015 0.871±0.0020 1.501±0.0009
Coupled (K=2) D1-F64-H8 0.349±0.0019 0.516±0.0011 0.700±0.0004 0.909±0.0023 1.144±0.0043 2.652±0.0109 0.432 0.057
Coupled (K=2) D4-F64-H8 0.313±0.0003 0.462±0.0003 0.622±0.0007 0.791±0.0024 0.982±0.0036 2.279±0.0059 0.483 0.201
Coupled (K=2) D8-F64-H8 0.318±0.0001 0.471±0.0016 0.633±0.0014 0.815±0.0041 1.035±0.0027 2.807±0.0118 0.550 0.394
Coupled (K=2) D8-F128-H8 0.563±0.0006 0.845±0.0007 1.136±0.0051 1.450±0.0052 1.791±0.0048 3.630±0.0141 0.994 1.552
Decoupled (K=2) (D1-F64-H8)×$\times 2$ 0.358±0.0006 0.532±0.0009 0.714±0.0027 0.919±0.0037 1.144±0.0057 2.613±0.0059 0.463 0.108
Decoupled (K=2) (D2-F64-H8)×$\times 2$ 0.340±0.0011 0.503±0.0006 0.676±0.0007 0.870±0.0015 1.080±0.0071 2.537±0.0049 0.496 0.200
Decoupled (K=2) (D4-F64-H8)×$\times 2$ 0.337±0.0006 0.503±0.0004 0.676±0.0034 0.866±0.0022 1.062±0.0047 2.409±0.0049 0.564 0.385
Table 4: Effect of Different Guidance. Our method outperforms CHOIS even without guidance. The feet-floor term improves physical realism, while the hand-object term enhances contact accuracy. Combining both achieves SOTA performance across all metrics.
Method Condition Matching Human Motion Interaction GT Difference
[…]
CHOIS (w/o gui) 1.86 5.79 3.05 3.39 0.57 3.63 0.54 0.42 0.61 16.05 25.34 12.36 0.99
CHOIS (w/ gui) 2.10 6.16 3.03 3.39 0.41 0.87 0.66 0.54 0.61 16.01 24.33 14.29 0.99
HOI-Dyn (w/o gui) 1.56 5.72 3.03 5.56 0.37 3.49 0.60 0.47 0.61 15.56 24.61 11.67 0.91
HOI-Dyn (feet-floor gui) 1.57 5.56 3.03 3.23 0.37 0.66 0.60 0.46 0.62 15.48 23.87 11.28 0.89
HOI-Dyn (hand-obj gui) 1.77 5.74 3.27 5.21 0.37 2.55 0.71 0.60 0.64 15.73 24.56 13.03 0.91
HOI-Dyn (w/ gui) 1.75 5.58 3.26 3.07 0.37 0.48 0.71 0.60 0.64 15.60 23.90 12.47 0.90
Figure 5: Frame-wise object point cloud loss. Our method HOI-Dyn (green line), consistently achieves lower loss than baselines without dynamics supervision (blue line). Notably, high loss peaks correspond to physically implausible HOI cases, which are effectively identified by our dynamics model.
5 Limitation and Conclusion
In this work, we emphasize the role of interaction dynamics in ensuring physical and causal consistency in HOI generation. As an early exploration, we propose a novel driver-responder framework and model interaction dynamics, integrating it with existing HOI motion diffusion techniques for more realistic interactions. However, our approach assumes rigid objects and uses the SMPL-X human model, which lack detailed hand representations, leading to minor inaccuracies in hand-object interactions, especially in rotations. For a more detailed discussion, see Appendix E.2. Looking ahead, we plan to incorporate more detailed object attributes and enhance human modeling to better capture fine-grained interactions [37]. We will also explore a tighter integration of HOI dependencies into generative models for improved interaction realism [38].
Appendix A Background of Classical Synchronization in Control Theory
To motivate our formulation, we revisit the classical driver-responder synchronization scheme from control theory, which forms the theoretical basis for modeling interaction dynamics in HOI generation.
[…]
where $x_{m},x_{s}\in\mathbb{R}^{n}$ are the state vectors of the Driver and Responder systems, respectively, $y_{i}\in\mathbb{R}^{p}$ ($i=m,s$) are the outputs, $F(\cdot)$ is a Lipschitz continuous nonlinear function, $B\in\mathbb{R}^{n\times r}$ is the constant input matrix, and $C\in\mathbb{R}^{p\times n}$ is the output matrix. The control input $u_{s}\in\mathbb{R}^{r}$ is given by
$u_{s}=K_{s}(y_{m}-y_{s})=K_{s}C(x_{m}-x_{s}),\quad K\in\mathbb{R}^{r\times n},$ (15)
where $K$ is the coupling strength. The synchronization objective is defined as
$\lim_{t\to\infty}\|x_{m}(t)-x_{s}(t)\|=0.$ (16)
Appendix B Details of HOI Motion and Condition Representation
Human Motion.
To represent the human motion $H$, we employ the SMPL-X parametric model, which reconstructs the 3D human mesh from pose and shape parameters. The pose is represented as a 204-dimensional vector, consisting of the following components: (i) a 3-dimensional translation vector that defines the joint position of the body, contributing a total of $24\times 3$ dimensions, and (ii) the rotations of 22 body joints, contributing $22\times 6$ dimensions to the pose vector. In addition to the pose parameters, the model includes a shape parameter vector $\boldsymbol{\beta}\in\mathbb{R}^{16}$, which controls the individual body shape.
Object Motion.
Object motion $O$ is characterized by two main components: the global 3D position and the relative rotation. The global position is represented by the centroid of the object, while the relative rotation, denoted as $\mathcal{R}_{\text{rel}}(t)$, describes the rotation at frame $t$ relative to the object’s geometry $P$. The object vertices at frame $t$ are given by $P_{t}=\mathcal{R}_{\text{rel}}(t)P$. Therefore, the object motion is described by $O\in\mathbb{R}^{T\times 12}$, where each frame consists of a translation relative to the centroid and a corresponding relative rotation.
Condition.
In our condition representation, we integrate contextual cues following the CHOIS framework. The text embedding is extracted using a pretrained CLIP text encoder. The object geometry is encoded via a MLP applied to its Basis Point Set (BPS), and the resulting features are broadcast to all frames. These BPS features are then concatenated with the initial states and waypoints (provided every 30 frames, where only the object’s $x$- and $y$-coordinates are available, with the $z$-axis omitted) and further combined with the text embedding to form the condition $\mathbf{c}$. In addition, due to the fact that we include a 4-dimensional binary contact indicator that specifies hand and feet contact states in HOI data representation, we therefore construct a masked motion representation to represent the initial states and waypoint constraints $\mathbf{m}\in\mathbb{R}^{T\times(12+204+4)}$, where the initial state includes both the human pose and object pose at the first frame.
Appendix C Theoretical and Empirical Justification of Interaction Dynamics Loss
C.1 Theoretical Justification via Residual Dynamics
Definition.
We define a learned dynamics model $\mathcal{D}$ that predicts the object’s relative motion based on the current scene and human motion:
${\Delta o}^{*}_{t\to t+k}\approx\mathcal{D}(s^{(t)},o^{(t)},\Delta h_{t\to t+k};\theta_{\mathcal{D}}),$ (17)
where $s^{(t)}$ is the scene context at time $t$, $o^{(t)}$ is the object state (e.g., pose), $\Delta h_{t\to t+k}$ is the future human motion, and $\theta_{\mathcal{D}}$ contains the model parameters.
To promote consistent physical behavior by the generated trajectories, we introduce an auxiliary loss that compares the residual dynamics behavior of generated and ground-truth sequences, as follows:
$\mathcal{L}_{\text{dyn}}=\left\|\delta_{\mathcal{D},\text{gen}}-\delta_{\mathcal{D},\text{gt}}\right\|^{2}$ (18)
with the residuals defined as
$\delta_{\mathcal{D},\text{gen}}$ $=\Phi\left(\mathcal{D}(\hat{s}^{(t)},\hat{o}^{(t)},{\Delta\hat{h}}_{t\rightarrow t+k}),{\Delta\hat{o}}_{t\to t+k}\right),$ (19)
$\delta_{\mathcal{D},\text{gt}}$ $=\Phi\left(\mathcal{D}(s^{(t)},o^{(t)},\Delta h_{t\rightarrow t+k}),\Delta o_{t\to t+k}\right),$
where $\Phi(\cdot,\cdot)$ denotes a point-wise distance metric (e.g., the $\ell_{1}$ distance between the point clouds decoded from object poses).
Bias Cancellation.
Although $\mathcal{D}$ may have approximation error, we show that comparing residuals (rather than raw predictions) reduces the effect of model bias. For theoretical clarity, we analyze in the prediction space without applying $\Phi$, and only consider $\Phi$ as a Lipschitz continuous final metric layer.
Let the true (unknown) dynamics function be $\mathcal{F}$, and define the model bias as
$b(x):=\mathcal{D}(x)-\mathcal{F}(x).$ (20)
Let $x_{\text{gen}}=(\hat{s}^{(t)},\hat{o}^{(t)},{\Delta\hat{h}}_{t\rightarrow t+k})$ and $x_{\text{gt}}=(s^{(t)},o^{(t)},\Delta h_{t\rightarrow t+k})$. Then the prediction-space residuals can be expressed as
$\delta_{\mathcal{D},\text{gen}}$ $=\mathcal{D}(x_{\text{gen}})-{\Delta\hat{o}}_{t\to t+k},$ (21)
$\delta_{\mathcal{D},\text{gt}}$ $=\mathcal{D}(x_{\text{gt}})-\Delta o_{t\to t+k}.$
Subsequently, the residual difference becomes
$\delta_{\mathcal{D},\text{gen}}-\delta_{\mathcal{D},\text{gt}}$ $=\left(\mathcal{D}(x_{\text{gen}})-\mathcal{D}(x_{\text{gt}})\right)+(\Delta o_{t\to t+k}-{\Delta\hat{o}}_{t\to t+k})$ (22)
$=\left[b(x_{\text{gen}})-b(x_{\text{gt}})\right]+\left[\mathcal{F}(x_{\text{gen}})-\mathcal{F}(x_{\text{gt}})\right]+(\Delta o-{\Delta\hat{o}}).$
Assuming that both $\mathcal{F}$ and $b$ are Lipschitz continuous with the Lipscthiz constants $L_{\mathcal{F}}$ and $L_{b}$, and that $\|x_{\text{gen}}-x_{\text{gt}}\|\leq\epsilon$, $\|{\Delta\hat{o}}-\Delta o\|\leq\delta$, then we obtain
$\left\|\delta_{\mathcal{D},\text{gen}}-\delta_{\mathcal{D},\text{gt}}\right\|\leq L_{b}\epsilon+L_{\mathcal{F}}\epsilon+\delta.$ (23)
This implies that
$\left\|\delta_{\mathcal{D},\text{gen}}-\delta_{\mathcal{D},\text{gt}}\right\|\to 0\quad\text{as}\quad\epsilon,\delta\to 0,$ (24)
meaning the residual discrepancy vanishes if the input conditions and object motion predictions are sufficiently close. Thus, the residual-based loss is less sensitive to the absolute accuracy of $\mathcal{D}$, and instead emphasizes consistency in behavior under the same model.
Connection to Training Loss.
While our analysis is carried out in the model output space, the actual training loss is applied using the decoded point cloud metric given by
$\Phi(x,y)=\left\|f_{\text{pc}}(x)-f_{\text{pc}}(y)\right\|_{1},$ (25)
where $f_{\text{pc}}(\cdot)$ denotes a deterministic decoder from object pose to point cloud. Since both $f_{\text{pc}}(\cdot)$ and $\|\cdot\|_{1}$ are Lipschitz continuous, the residual error behavior remains consistent under this transformation.
C.2 Empirical Evidence from Motion Decomposition
Setup.
To validate the approximation $\mathbb{E}_{t}[\delta_{\mathcal{D},\text{gen}}-\delta_{\mathcal{D},\text{gt}}]\approx 0$, we compute the residual difference
$\Delta^{(n)}(t)=\delta_{\mathcal{D},\text{gen}}^{(n)}(t)-\delta_{\mathcal{D},\text{gt}}^{(n)}(t)$ (26)
for each sequence $n\in\{1,2,\dots,N\}$ and time step $t\in\{1,2,\dots,T\}$. We then compute the aggregate statistics over all $NT$ samples, as follows:
$\bar{\Delta}=\frac{1}{NT}\sum_{n=1}^{N}\sum_{t=1}^{T}\Delta^{(n)}(t),\quad\text{var}(\Delta)={\frac{1}{NT}\sum_{n=1}^{N}\sum_{t=1}^{T}\left(\Delta^{(n)}(t)-\bar{\Delta}\right)^{2}}$ (27)
These metrics reflect the average residual discrepancy and its variability across all sequences and frames.
Table 5: Comparison of Statistical Measures.
Statistic Mean Variance
Ground Truth (gt) 0.3130 0.1190
Generated (gen) 0.3147 0.0950
Difference (gen - gt) 0.0016 0.1561
Figure 6: Statistical and distributional comparison of interaction dynamics errors for generated and ground-truth object motion. The mean and variance of residuals are nearly identical for gen and gt, while their difference is centered near zero with low variance. Probability density plots further confirm the alignment of gen and gt distributions, and support the theoretical claim of modeling bias cancellation.
Results.
As shown in Table 5, the mean of the residual difference is nearly zero ($0.0016$), confirming that the residual signal has no significant bias across time steps or sequences, consistent with our theoretical approximation $\mathbb{E}_{t}[\delta_{\mathcal{D},\text{gen}}-\delta_{\mathcal{D},\text{gt}}]\approx 0$.
We further visualize the typical temporal patterns in Fig. 6, where the residual differences fluctuate randomly around zero, with no observable drift or systematic deviation. These empirical observations validate the core assumption that residual-based comparison cancels out the model bias when $\mathcal{D}$ is applied consistently to both generated and ground-truth trajectories.
Together with our theoretical insights, these results support the following: (i) The residual-based loss introduces natural bias cancellation, making the learning robust to approximation errors in $\mathcal{D}$. (ii) This enables the use of $\mathcal{D}$ as an auxiliary evaluator without requiring high-fidelity absolute predictions. (iii) The formulation enforces consistency in human-object dynamics, thereby improving the realism and coherence of generated motion. This bias-invariant design is a key strength of our auxiliary loss and contributes to the stability, generalizability, and interpretability of the learned policy.
Appendix D Additional Qualitative Results of HOI Generation
D.1 Generated HOI Sequences
We provide additional qualitative comparisons to further evaluate the effects of modeling interaction dynamics.
In Fig. 7, we assess the Action–Interaction–Response loop. The results show that incorporating interaction dynamics constraints enables the model to better understand how objects should respond to both human actions and the contact context. This suggests that modeling such dynamics improves physical plausibility and contextual coherence in human–object interactions.
In Fig. 8, we further investigate sequence-level alignment. We observe that explicitly enforcing interaction dynamics not only enhances the consistency of human–object contact but also improves the overall motion quality. These findings align with our claim: “Contact is implicitly governed by the dynamics—no need to explicitly model it. If there is no contact, there is no response; if contact occurs, the object’s response is naturally determined by the interaction dynamics.”
Moreover, by restricting the object from moving in physically implausible ways, our method encourages the human to take more active and reasonable actions. This results in motion that is not only more coordinated but also better aligned with the intended conditions or text prompts—addressing a key limitation in prior works where human and object often move independently without dynamic coherence.
Figure 7: Qualitative comparison in Action–Interaction–Response.
Figure 8: Qualitative comparison in Sequence-level alignment.
D.2 Applications in 3D Scene
We demonstrate the practical applicability of our method by synthesizing dynamic human–object interactions within realistic 3D environments, conditioned on natural language descriptions.
Specifically, we use 3D scenes from the Replica dataset and manually define interaction instructions such as “pull the floor lamp and move it next to the sofa”. The pipeline involves (i) parsing the textual description to identify the intended interaction and the associated object(s), (ii) specifying the target object’s initial placement within the scene, and (iii) planning a collision-free navigation trajectory using Habitat. The resulting path is then provided to our HOI generation model, which produces coherent and physically plausible motion sequences aligned with the instruction.
As illustrated in Fig. 9, our model successfully generates realistic agent behaviors within the scene, demonstrating interaction fidelity and environmental awareness. This technique has broad potential across domains such as animation production, interactive scene synthesis, and virtual reality. Moreover, it provides promising avenues for applications in embodied intelligence, sim-to-real transfer, imitation learning, and human-centered robotics.
Figure 9: HOI generation in a realistic 3D scene. The virtual agent interacts with a floor lamp as instructed, producing physically consistent motion with respect to the environment.
Appendix E Network and Additional Results of Interaction Dynamics
E.1 Model Architecture
Our dynamics model $\mathcal{D}$ is a lightweight predictor designed to infer object motion induced by human actions and interaction context. It takes as input the object-centric BPS point cloud, human joint relative motions, and contact information, and outputs the relative object motion (rotation + translation). The architecture consists of three main components:
(i) BPS Encoder. A two-layer multilayer perceptron (MLP) encodes the input object BPS ($1024\times 3$) into a compact feature vector. This captures the object’s shape and spatial context.
(ii) Human Motion Module. A Mini Transformer models the temporal interaction dynamics between the human and the object. It first encodes the joint-level motion features (3D relative positions and rotations) through a MLP, then injects condition information—concatenation of object motion, contact state, and BPS feature—via conditional encoding. A shallow transformer with learnable positional embeddings is applied, followed by attention pooling to obtain a global human motion feature.
(iii) Motion Decoder. A MLP maps the fused interaction feature to the target object motion vector ($3+9$ dimensions), representing translation and flattened rotation. The output can be optionally projected to $\text{SO}(3)$ to ensure a valid rotation.
E.2 One-Step and Auto-Regressive Prediction Results
We evaluate the predictive ability of our interaction dynamics model from two complementary perspectives: one-step prediction and auto-regressive prediction.
In the one-step setting, the model is given the current scene context and ground-truth human motion to predict the object state at the next frame ($T=1$). This setting directly evaluates the local dynamics response. As shown in Fig. 13, our model achieves high accuracy in object translation, with very small deviations from ground truth. However, the rotation prediction still shows room for improvement. We attribute this to the dataset lacking hand pose details, which limits the inference of fine-grained object rotations from body joints alone. Nevertheless, the predicted rotations remain broadly consistent with the ground truth.
In the auto-regressive setting, we recursively feed the predicted object state back into the model to generate long-term future motion. This setting reveals how prediction errors accumulate. As illustrated in Fig. 13, even though our model is only trained with one- and two-step supervision, it demonstrates impressive generalization in the auto-regressive regime. The predicted object translations stay closely aligned with the ground truth over time. Although rotation errors do increase due to compounding uncertainty, they remain stable and plausible.
Importantly, our method relies solely on the one-step supervision during training, which avoids the inefficiency and instability of full auto-regressive rollout training. This design enables fast training while already delivering strong performance. The model’s ability to generalize well in the auto-regressive case is a bonus, not a requirement. Hence, the proposed design achieves an excellent balance between efficiency and long-horizon performance — a highly cost-effective formulation.
E.3 Comparison on Interaction Dynamics
We further analyze the quality of interaction dynamics generated by CHOIS and our proposed HOI-Dyn. As shown in Fig. 13 and Fig. 13, we employ the predicted human-object sequences and compare their induced object motion against the reference dynamics. Here, the reference is not the ground truth in the test set , but the object’s relative motion in the synthetic sequences used as initial conditions for prediction. Since motion synthesis is inherently multimodal, we focus on whether the predicted motion is physically and contextually plausible.
We observe that HOI-Dyn produces significantly more consistent and coherent object motion. In contrast, CHOIS often yields noisy or unstable object transitions, with noticeable discontinuities or implausible drifts over time.This comparison highlights the advantage of our synchronized control formulation and the dynamics-guided residual loss.
Figure 10: One-step performance.
Figure 11: Auto-regressive performance.
Figure 12: CHOIS.
Figure 13: HOI-Dyn.
E.4 Joint-Level Attention Visualization
In our proposed driver-responder synchronized control framework, we posit that once contact is established, the object’s motion should predominantly follow the human motion. Based on the SMPL representation, human motion is determined by the transitions and rotations of body joints. We thus interpret object motion as a result of the aggregated influence of all human joints.
Our MiniTransformer module explicitly models joint-wise contributions to object dynamics by applying attention over all 24 SMPL joints. Specifically, given joint features $\mathbf{x}_{j}\in\mathbb{R}^{d}$ for joint $j\in\{1,\dots,24\}$, and a learned attention weight $w_{j}$, we compute the global representation as:
[…]
where the attention weights $w_{j}\in[0,1]$ are computed via a softmax layer and satisfy $\sum_{j}w_{j}=1$. This formulation allows the model to selectively focus on the most relevant joints under different interaction conditions, enabling interpretable and context-aware reasoning of human-object dynamics.
We visualize the joint-level attention maps for typical cases in Fig. 15 and Fig. 15. For the first example in Fig. 15, the person bends down to pick up a box. During this full-contact phase, the object is entirely under human control. The attention map highlights not only the hands but also the feet, elbows, and chest — suggesting a whole-body coordination that is both intuitive and desirable. This supports the idea of a “global response” where object motion is governed by distributed joint influence.
For another example in Fig. 15, where the person is walking while carrying the box, the attention is more concentrated on the hands and feet. This makes intuitive sense: the hands are directly manipulating the object (contributing to both translation and rotation), while the feet govern the global transition through locomotion. Across multiple sequences, we also observe consistent attention on the hips and lower limbs, suggesting their role as stable references for interpreting whole-body movement.
Figure 14: Pick up a box.
Figure 15: Walk forward while holding the box.
Appendix F Details of Classifier-based Guidance
While the use of guidance modules (e.g., contact-based constraints) can lead to targeted improvements in motion quality, all comparisons with state-of-the-art baselines are conducted under identical guidance settings. This ensures a fair evaluation, as our proposed method HOI-Dyn focuses on internally controlling and optimizing the interaction dynamics during the generation process. Therefore, the guidance configuration is kept fully consistent across all methods and does not introduce any bias in favor of HOI-Dyn.
F.1 Feet-Floor Contact Guidance
To encourage physical plausibility during locomotion and standing still, we introduce a contact guidance loss that penalizes unnatural foot height above the ground, consistent with the CHOIS setting. Specifically, we extract the 3D positions of the left and right toe joints, and compute the support foot height as:
$h_{\text{support}}=\min\left(h_{\text{left-toe}},\ h_{\text{right-toe}}\right),$ (29)
where $h\in\mathbb{R}^{B\times T\times 1}$ represents the height (Z-axis) across batch $B$ and time $T$.
We enforce that the support foot remains near the ground $\alpha$ (set at 2 cm) via an MSE loss:
$\mathcal{L}_{\text{feet-floor}}=\frac{1}{BT}\sum_{b=1}^{B}\sum_{t=1}^{T}\left(h_{\text{support}}^{(b,t)}-\alpha\right)^{2}.$ (30)
This loss improves contact realism by promoting plausible foot-ground interaction.
F.2 Hand-Object Contact Guidance
To enhance the realism and stability of hand-object interactions, we use the multi-term auxiliary loss that encourages physically plausible contact and temporal coherence.
  • Contact Loss: For each timestep, we extract the 3D positions of the left and right palm joints and compute their closest distances to the object mesh as follows:
    $\mathcal{L}_{\text{contact}}=\frac{1}{BT}\sum_{b=1}^{B}\sum_{t=1}^{T}\sum_{h\in\{\text{L,R}\}}\left[\max(\|\mathbf{p}_{h}^{(b,t)}-\mathcal{M}^{(b,t)}\|-\delta,0)\cdot\phi_{\text{contact}}^{(b,t,h)}\right],$ (31)
    where $\mathcal{M}^{(b,t)}$ is the object mesh at frame $t$, $\mathbf{p}_{h}^{(b,t)}$ is the palm position, $\delta=0.02$ is the contact threshold, and $\phi_{\text{contact}}$ is the predicted binary contact indicator.
  • Temporal Consistency Loss: For contact frames, we enforce that the palm positions (in the object coordinate frame) remain temporally smooth and aligned by
    $\mathcal{L}_{\text{consistency}}=2-\frac{1}{T^{2}}\sum_{i,j}\left[\cos(\theta^{\text{L}}_{i,j})\cdot M_{i,j}^{\text{L}}+\cos(\theta^{\text{R}}_{i,j})\cdot M_{i,j}^{\text{R}}\right],$ (32)
    where $\theta_{i,j}$ is the angle between the relative palm directions at frames $i$ and $j$, and $M^{\text{L/R}}_{i,j}$ are symmetric masks marking the frames where contact is active.
  • Floor Penetration Penalty: We suppress object vertices penetrating below the ground plane $(z<0)$ through
    $\mathcal{L}_{\text{floor}}=\frac{1}{N}\sum_{n=1}^{N}\max(-z_{n},0).$ (33)
The final guidance loss is given by
$\mathcal{L}_{\text{interaction}}=\mathcal{L}_{\text{contact}}+\mathcal{L}_{\text{consistency}}+\lambda\cdot\mathcal{L}_{\text{floor}},\quad\text{with}\quad\lambda=100.$ (34)
This auxiliary loss encourages accurate and stable contact modeling.
Appendix G Additional Results after Submission
Table 6: Residual vs Non-Residual Loss.
Method Condition Matching Human Motion Interaction GT Difference
[…]
w/o gui
CHOIS 1.86 5.79 3.05 3.39 0.57 3.63 0.54 0.42 0.61 16.05 25.34 12.36 0.99
HOI-Dyn (Res Loss) 1.56 5.72 3.03 5.56 0.37 3.49 0.60 0.47 0.61 15.56 24.61 11.67 0.91
HOI-Dyn (NonRes Loss) 1.56 5.24 3.03 5.33 0.39 xxx 0.59 0.47 0.71 15.63 25.10 11.86 0.89
w/ gui
CHOIS 2.10 6.16 3.03 3.39 0.41 0.87 0.66 0.54 0.61 16.01 24.33 14.29 0.99
HOI-Dyn (Res Loss) 1.75 5.58 3.26 3.07 0.37 0.48 0.71 0.60 0.64 15.60 23.90 12.47 0.90
HOI-Dyn (NonRes Loss) 1.69 5.33 3.07 2.98 0.40 xxx 0.73 0.64 0.70 15.60 24.31 12.79 0.88
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Claims we made accurately reflect the paper’s contributions and scope.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The interaction dynamics model may not capture object rotations with high precision, mainly due to limited data fidelity or missing fine-grained physical cues. However, our residual-based interaction dynamics loss is specifically designed to mitigate such inaccuracies by focusing on differential consistency rather than absolute correctness.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: We provide a comprehensive justification for the robustness of our residual-based interaction dynamics loss to imperfect model learning, grounded in clear assumptions, theoretical analysis, and empirical evidence.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Experimental setups are described in detail.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [No]
  24. Justification: Will release our code after acceptation, or we can provide the code if any reviewers are interested in it during the review process. Anyway, the details about the code have already described in the paper.
  25. Guidelines: The code will be released upon acceptance. We are also willing to share it with reviewers during the review process upon request. All necessary implementation details are included in the paper.
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We have described all the details about of experiments.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We have conducted statistical tests in our experiments, error bars are plotted in figures.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: All computational resources used in our experiments are reported in detail in the experimental setup. Notably, our method is highly efficient: all experiments were conducted on a single NVIDIA RTX A4500 GPU with 20GB memory. We also provide an analysis of model parameters and inference cost, demonstrating that our interaction dynamics model is lightweight.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [N/A]
  44. Justification: Not applicable.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [No]
  49. Justification: Our algorithm has no potential negative social impacts.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [No]
  54. Justification: Code will be released after acceptation, it would be open access, no safeguards are required.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We have cited the existing assets we used in our paper.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [No]
  64. Justification: We did not introduce any new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: We do not have any experiments or research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: We do not have any experiments or research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [No]
  79. Justification: No large language models (LLMs) were involved in the design, development, or evaluation of the core methods presented in this paper. LLMs were only used as a writing assistant to improve the clarity and fluency of the text. As such, their usage does not impact the scientific rigor, methodology, or originality of the work.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.